# Token Interdependency Parsing (Tipping) - Fast and Accurate Log Parsing

### v1

@@ 4.3.1 Experiment depiction
LogLead [^6] https://github.com/EvoTestOps/LogLead is a tool designed to benchmark log representations and anomaly detection algorithms, such as log parsing results [35]. It features integrations with many public datasets and log parsers. LogLead is built with efficiency in mind, achieved by using the Polars [^7] https://github.com/pola-rs/polars library, a notably faster alternative to Pandas [^8] https://github.com/pandas-dev/pandas. As LogLead integrates existing datasets and parsers, it offers an environment to evaluate the computational efficiency of different log parsers over different datasets. Unlike the experiments in the previous sections, no dataset-specific or parser-specific optimizations were made, and as such, are not available on the LogLead platform.
The scripts we used for our LogLead experiments are available [^9] https://github.com/EvoTestOps/LogLead/tree/main/demo/parser_benchmark. For the purposes of this experiment, we first loaded and preprocessed log data from nine different systems from past studies. This preprocessing involved loading data into Polars data frames and normalizing it using a standard regular expression in the LogLead Enhancer class. The datasets included BGL, Liberty, Spirit, and Thunderbird, originating from real-world high-performance computing usage [36]. The Hadoop[37] and HDFS[38] datasets contain logs from experiments conducted on distributed file systems. Additionally, Trainticket [39] and Web-shop [^10] https://github.com/GoogleCloudPlatform/microservices-demo datasets represent experiment logs from modern microservice-based systems [40]. Finally, the Android2 dataset, not to be confused with the Android dataset, originates from long-lasting Android reliability testing runs [41].
@@
Parser implementations for AEL, Brain, and IpLom were taken from [24]. For Spell [^11] https://github.com/bave/pyspell and LenMa [^12] https://github.com/keiichishima/templateminer, we used the implementations available in the footnotes. For Drain, we used the so-called Drain3131313https://github.com/logpai/Drain3/ implementation. In the Drain configuration file, we disabled Drain’s internal log message normalization via regular expressions and disabled profiling to ensure maximal parsing speed. For other parsers, no such settings were available.
As our purpose was to investigate the benefits of parallelization, we decided to implement a paralyzation version of the IpLom parser. We named it PlIpLom, Pl being short for Polars. We followed the original IpLom paper and the Python-based source code implementation from [24], but whenever we encountered a loop, we replaced it with native Polars operations if possible. Native Polars operations are known for their efficiency.
@@ 4.3.2 Experiment results
During our experiments, we set a run-time limit of 10 minutes (600 seconds for each parser on each dataset). Accordingly, if a run exceeded 10 minutes, we allowed it to finish, but no further runs were executed. The remaining results are marked as ”Out of Time Limit” (OTL) in the tables from VII to XXI. We also mark crashed runs as ”Did Not Finish” (DNF). We conducted a re-run for all crashes; however, if the re-run also failed to complete, we marked it as DNF.
@@
We have two small log datasets. The results for the Hadoop dataset (177,592 log lines) on the local machine (LM) and virtual machine (VM) are shown in Tables VII and VIII, respectively. Tipping, Drain, and PlIplom are the three fastest parsers in both the VM and LM settings. Similar to Hadoop, TrainTicket is a small dataset (272,270 log lines), and its results are documented in Tables IX and X. In the TrainTicket dataset, Drain is the fastest on the LM, followed by Tipping and Iplom. However, the performance ranking on the VM changes, with Tipping being the fastest, followed by Drain and Iplom.
TABLE VII: Parsing times (in seconds) on the local machine using Hadoop dataset for various parsers across different dataset sizes, ranging from 1/1024 to the full dataset.
Method 1/1024 1/512 1/256 1/128 1/64 1/32 1/16 1/8 1/4 1/2 1/1
AEL 0.04 0.02 0.04 0.07 0.12 0.20 0.46 0.79 1.62 3.24 6.41
Brain 0.03 0.02 0.03 0.07 0.13 0.29 0.64 1.15 2.43 5.07 9.81
Drain 0.01 0.00 0.01 0.01 0.02 0.03 0.06 0.12 0.23 0.47 0.93
Iplom 0.01 0.01 0.01 0.02 0.04 0.07 0.13 0.30 0.66 1.23 2.29
Lenma 0.05 0.04 0.09 0.20 0.41 0.86 1.49 3.41 7.16 15.73 29.54
Pliplom 0.16 0.23 0.30 0.38 0.45 0.52 0.63 0.73 0.89 1.10 1.66
Spell 0.02 0.02 0.06 0.14 0.33 0.82 1.84 4.50 10.73 23.04 47.52
Tipping 0.00 0.00 0.00 0.00 0.01 0.01 0.02 0.04 0.07 0.11 0.22
TABLE VIII: Parsing times (in seconds) on the virtual machine using Hadoop dataset for various parsers across different dataset sizes, ranging from 1/1024 to the full dataset.
Method 1/1024 1/512 1/256 1/128 1/64 1/32 1/16 1/8 1/4 1/2 1/1
AEL 0.07 0.03 0.06 0.10 0.22 0.43 0.80 1.53 3.29 6.77 13.29
Brain 0.04 0.04 0.07 0.15 0.41 0.51 1.21 2.17 4.96 11.15 21.18
Drain 0.02 0.01 0.01 0.02 0.04 0.08 0.15 0.31 0.56 1.18 2.22
Iplom 0.04 0.02 0.03 0.06 0.08 0.15 0.27 0.56 1.01 2.17 5.01
Lenma 0.06 0.10 0.21 0.45 0.94 1.99 4.35 9.27 19.05 40.97 86.78
PlIplom 0.20 0.30 0.29 0.45 0.53 0.66 0.82 1.02 1.38 1.86 2.80
Spell 0.04 0.05 0.14 0.34 0.91 2.07 4.54 10.08 21.68 49.11 105.34
Tipping 0.03 0.01 0.01 0.01 0.01 0.02 0.03 0.05 0.11 0.20 0.36
TABLE IX: Parsing times (in seconds) on the local machine using Nezha-TrainTicket dataset for various parsers across different dataset sizes, ranging from 1/1024 to the full dataset.
Method 1/1024 1/512 1/256 1/128 1/64 1/32 1/16 1/8 1/4 1/2 1/1
AEL 0.35 0.25 0.43 0.39 0.39 0.89 1.50 2.83 5.69 9.90 16.86
Brain 0.04 0.07 0.10 0.19 0.40 0.83 1.62 3.32 6.19 22.62 32.58
Drain 0.00 0.01 0.01 0.01 0.03 0.05 0.09 0.17 0.33 0.89 1.74
Iplom 0.01 0.02 0.03 0.05 0.09 0.19 0.38 0.72 1.42 3.36 6.28
Lenma 0.03 0.07 0.15 0.35 0.72 1.59 3.38 7.83 18.46 48.39 122.89
Pliplom 0.72 0.76 0.92 1.09 1.33 1.76 2.48 4.40 5.13 7.88 9.69
Spell 0.05 0.13 0.20 0.49 1.09 2.00 4.15 8.42 18.18 40.46 78.96
Tipping 0.01 0.01 0.01 0.02 0.03 0.03 0.06 0.12 0.43 1.05 2.01
TABLE X: Parsing times (in seconds) on the virtual machine using Nezha-TrainTicket dataset for various parsers across different dataset sizes, ranging from 1/1024 to the full dataset.
Method 1/1024 1/512 1/256 1/128 1/64 1/32 1/16 1/8 1/4 1/2 1/1
AEL 0.39 0.44 0.68 0.72 0.95 1.80 2.77 5.63 11.02 21.25 37.60
Brain 0.11 0.12 0.16 0.33 1.30 2.06 3.12 7.49 15.18 30.17 63.20
Drain 0.05 0.01 0.02 0.04 0.07 0.12 0.24 0.46 0.89 1.83 3.70
Iplom 0.07 0.05 0.06 0.11 0.22 0.42 0.82 1.69 2.98 6.43 13.55
Lenma 0.17 0.23 0.46 1.13 1.98 4.29 9.25 21.20 51.90 135.89 349.02
Pliplom 0.89 1.22 1.30 1.98 2.15 2.91 4.17 5.83 8.42 10.35 14.72
Spell 0.22 0.31 0.45 1.05 2.13 4.03 8.38 17.54 40.17 83.17 164.07
Tipping 0.02 0.02 0.01 0.02 0.02 0.04 0.09 0.18 0.34 0.62 1.11
Next, we have four medium-sized datasets.
@@
Web-shop, with 3,958,203 log lines, presents its results for the local machine (LM) and virtual machine (VM) in Tables XI and XII, respectively. In the LM setting, the top three parsers are PlIplom, Tipping, and Drain, whereas in the VM, the order is Tipping, PlIplom, and Drain. Here, we begin to observe meaningful differences in parsing speed, with the difference between the first and third parsers approximately 9x in LM and 8x in VM.
@@
The results for the BGL dataset, containing 4,747,963 log lines, are displayed in Tables XIII and XIV. In both computing environments, the top three parsers, in order, are Tipping, PlIplom, and Drain.
@@
The Android2 dataset, with 5,203,599 log lines, shows results in Tables XV and XVI. Tipping is the fastest parser in both the VM and LM, while the second and third places alternate between Drain and PlIplom in the VM and LM, respectively.
@@
Finally, the HDFS dataset, which consists of 11,175,629 log lines, is presented in Tables XVII and XVIII. In the LM, PlIplom is the fastest, followed by Tipping and Drain, while in the VM, Tipping leads, followed by PlIplom and Drain.
TABLE XI: Parsing times (in seconds) on the local machine using Nezha-Shop dataset for various parsers across different dataset sizes, ranging from 1/1024 to the full dataset.
Method 1/1024 1/512 1/256 1/128 1/64 1/32 1/16 1/8 1/4 1/2 1/1
AEL 0.17 0.27 0.48 0.92 2.10 3.60 7.06 13.38 28.69 58.08 119.08
Brain 0.16 0.33 0.71 1.35 2.70 6.03 12.12 23.45 46.38 95.54 191.24
Drain 0.02 0.03 0.06 0.14 0.31 0.57 1.38 2.79 5.71 9.90 25.61
Iplom 0.04 0.07 0.14 0.28 0.57 1.40 3.07 6.06 10.61 19.46 39.52
Lenma 1.34 3.06 6.70 12.60 26.86 57.47 126.78 244.08 470.02 1660.49 OTL
Pliplom 0.76 1.33 1.06 1.20 1.44 1.47 1.59 1.89 2.39 3.07 4.73
Spell 0.27 0.61 1.26 3.25 6.26 14.33 32.68 75.75 164.32 364.17 1556.72
Tipping 0.01 0.02 0.03 0.07 0.09 0.24 0.52 0.99 2.07 3.62 7.69
TABLE XII: Parsing times (in seconds) on the virtual machine using Nezha-Shop dataset for various parsers across different dataset sizes, ranging from 1/1024 to the full dataset.
Method 1/1024 1/512 1/256 1/128 1/64 1/32 1/16 1/8 1/4 1/2 1/1
AEL 0.32 0.47 1.01 1.80 3.73 7.83 16.45 31.66 68.67 139.55 291.96
Brain 0.37 0.57 1.01 2.08 4.18 9.65 20.32 40.82 91.96 181.47 357.70
Drain 0.15 0.09 0.17 0.33 0.71 1.35 2.68 5.67 10.62 20.02 40.52
Iplom 0.25 0.13 0.25 0.51 1.03 2.44 4.47 11.43 28.73 54.01 106.90
Lenma 2.63 5.58 11.13 23.70 50.15 103.71 225.87 511.49 1183.57 OTL OTL
Pliplom 0.61 0.53 0.56 0.99 1.24 1.11 1.58 2.09 3.84 6.53 11.52
Spell 0.57 1.16 2.76 6.86 15.81 34.60 79.54 184.76 417.56 886.77 OTL
Tipping 0.08 0.02 0.04 0.08 0.14 0.25 0.43 0.92 1.72 3.12 5.89
TABLE XIII: Parsing times (in seconds) on the local machine using BGL dataset for various parsers across different dataset sizes, ranging from 1/1024 to the full dataset.
Method 1/1024 1/512 1/256 1/128 1/64 1/32 1/16 1/8 1/4 1/2 1/1
AEL 0.20 0.37 0.66 1.44 3.38 5.72 11.78 23.58 48.73 96.24 348.72
Brain 0.16 0.31 0.66 1.25 3.16 5.38 11.75 22.68 44.08 90.17 169.00
Drain 0.03 0.05 0.09 0.17 0.33 0.75 1.36 2.71 5.51 12.68 21.14
Iplom 0.06 0.09 0.19 0.37 0.76 1.39 3.46 7.99 14.61 30.28 43.58
Lenma 6.80 15.99 39.01 96.63 223.21 611.27 OTL OTL OTL OTL OTL
Pliplom 1.05 1.32 1.58 1.89 2.33 2.84 3.14 4.11 5.12 7.69 12.18
Spell 0.45 1.19 2.96 5.81 13.39 29.93 68.99 168.33 424.29 975.84 OTL
Tipping 0.03 0.02 0.04 0.07 0.15 0.42 0.69 1.44 2.78 5.87 8.86
@@
TABLE XIV: Parsing times (in seconds) on the virtual machine using BGL dataset for various parsers across different dataset sizes, ranging from 1/1024 to the full dataset.
Method 1/1024 1/512 1/256 1/128 1/64 1/32 1/16 1/8 1/4 1/2 1/1
AEL 0.40 0.61 1.23 2.40 4.50 10.01 20.36 42.46 87.72 177.31 639.22
Brain 0.36 0.64 1.33 2.36 4.36 11.97 26.45 53.28 101.79 186.76 354.50
Drain 0.06 0.08 0.16 0.34 0.67 1.35 2.87 5.61 11.81 24.08 44.20
Iplom 0.18 0.16 0.32 0.62 1.11 2.12 6.19 12.56 28.00 57.47 100.47
Lenma 12.42 36.92 98.55 226.25 522.69 1091.13 OTL OTL OTL OTL OTL
Pliplom 0.55 0.55 0.75 0.86 1.22 1.63 2.22 2.99 4.63 8.78 15.03
Spell 0.75 1.59 3.84 8.56 20.99 48.27 110.90 254.86 608.05 OTL OTL
Tipping 0.06 0.03 0.05 0.08 0.15 0.27 0.49 1.00 1.95 3.90 7.35
TABLE XV: Parsing times (in seconds) on the local machine using Android dataset for various parsers across different dataset sizes, ranging from 1/1024 to the full dataset.
Method 1/1024 1/512 1/256 1/128 1/64 1/32 1/16 1/8 1/4 1/2 1/1
AEL 0.19 0.35 0.74 1.37 2.71 6.63 13.53 40.17 142.04 521.12 1657.02
Brain 0.19 0.35 0.79 1.45 3.21 7.36 14.81 27.96 53.83 109.59 210.54
Drain 0.04 0.05 0.09 0.18 0.51 0.95 2.02 4.65 7.96 15.79 30.56
Iplom 0.06 0.09 0.19 0.41 0.80 1.75 4.50 8.35 16.41 31.51 57.86
Lenma 2.12 5.84 14.26 31.20 73.15 174.46 386.23 803.56 OTL OTL OTL
Pliplom 1.88 3.12 4.20 5.93 7.78 10.62 12.56 16.02 20.70 26.13 33.04
Spell 2.66 6.18 23.36 56.15 150.84 416.88 1096.78 OTL OTL OTL OTL
Tipping 0.04 0.02 0.05 0.09 0.16 0.32 0.65 1.37 2.73 5.19 9.44
TABLE XVI: Parsing times (in seconds) on the virtual machine using Android dataset for various parsers across different dataset sizes, ranging from 1/1024 to the full dataset.
Method 1/1024 1/512 1/256 1/128 1/64 1/32 1/16 1/8 1/4 1/2 1/1
AEL 0.58 0.66 1.81 3.20 6.28 12.39 29.96 79.40 230.08 1268.74 OTL
Brain 0.48 0.70 1.39 2.72 6.87 13.14 29.47 61.62 118.08 243.72 468.88
Drain 0.12 0.12 0.24 0.43 0.84 1.74 4.08 7.24 14.90 31.40 60.84
Fiplom 14.92 21.72 32.20 43.48 55.78 65.95 77.89 98.67 118.54 145.41 170.02
Iplom 0.18 0.17 0.44 0.85 1.56 3.49 7.75 15.55 32.01 62.45 119.40
Lenma 3.67 8.15 19.64 47.01 112.64 264.75 606.25 OTL OTL OTL OTL
Pliplom 1.64 2.65 3.51 4.88 6.15 7.76 10.46 13.12 18.61 24.98 37.50
Spell 3.35 10.20 27.17 72.79 201.45 514.25 1313.77 OTL OTL OTL OTL
Tipping 0.15 0.04 0.06 0.12 0.19 0.34 0.57 1.16 2.15 4.49 8.82
TABLE XVII: Parsing times (in seconds) on the local machine using HDFS dataset for various parsers across different dataset sizes, ranging from 1/1024 to the full dataset.
Method 1/1024 1/512 1/256 1/128 1/64 1/32 1/16 1/8 1/4 1/2 1/1
AEL 0.43 0.65 1.43 3.43 7.13 10.94 20.66 40.94 80.88 165.51 338.08
Brain 0.76 1.13 1.89 4.64 6.63 19.56 34.24 63.73 128.22 255.77 DNF
Drain 0.08 0.22 0.21 0.42 0.85 1.80 4.21 9.11 14.81 29.80 76.20
Iplom 0.23 0.29 0.87 1.39 3.11 4.98 10.23 19.07 32.74 64.23 119.42
Lenma 3.24 6.56 11.94 19.69 37.69 75.23 151.22 301.56 612.81 OTL OTL
Pliplom 0.13 0.10 0.15 0.19 0.29 0.50 0.87 1.60 3.04 5.05 8.89
Spell 0.50 1.01 2.08 4.54 9.39 19.95 43.19 93.35 204.90 410.00 821.56
Tipping 0.05 0.05 0.06 0.09 0.17 0.39 0.95 1.98 3.99 9.38 16.44
TABLE XVIII: Parsing times (in seconds) on the virtual machine using HDFS dataset for various parsers across different dataset sizes, ranging from 1/1024 to the full dataset.
Method 1/1024 1/512 1/256 1/128 1/64 1/32 1/16 1/8 1/4 1/2 1/1
AEL 1.05 1.55 3.05 5.93 11.08 22.50 45.16 90.46 182.24 358.16 727.55
Brain 0.99 1.51 3.32 6.08 12.31 26.61 62.69 122.65 246.63 478.24 1016.13
Drain 0.17 0.27 0.50 1.01 1.65 3.39 6.47 13.33 28.02 58.52 128.90
Iplom 0.48 0.52 0.97 1.74 3.43 7.17 14.89 37.84 70.83 142.33 256.24
Lenma 5.10 9.83 20.42 39.47 82.55 168.56 335.38 687.35 OTL OTL OTL
Pliplom 0.26 0.19 0.26 0.43 0.58 1.07 1.95 4.30 8.62 16.33 31.21
Spell 0.94 1.47 3.17 6.67 15.57 32.42 69.31 144.91 328.61 697.07 OTL
Tipping 0.22 0.05 0.10 0.19 0.34 0.64 1.06 2.26 4.71 8.84 16.60
Finally, we have three large datasets executed only in the virtual machine (VM), as the local machine’s memory size was not sufficient. In the VM, none of the parsing executions resulted in fully parsed datasets, as all of the parsers either exceeded the time limit or ran out of memory.
The Thunderbird dataset, with 211,212,192 log lines, shows parsing results in Table XIX. The fastest parsers, in order, were Tipping, Iplom, and Drain. This order is consistent across the remaining two large datasets. The Liberty and Spirit datasets’ results are shown in Tables XXI and XX, respectively. Compared to its competitors, the only drawback noted for Tipping is its inability to parse half of the Liberty data, a task that PlIplom manages to accomplish.
TABLE XIX: Parsing times (in seconds) on the virtual machine using Thunderbird dataset for various parsers across different dataset sizes, ranging from 1/1024 to 1/2 of the dataset.
@@
TABLE XX: Parsing times (in seconds) on the virtual machine using Spirit dataset for various parsers across different dataset sizes, ranging from 1/1024 to 1/4 of the dataset.
@@
TABLE XXI: Parsing times (in seconds) on the virtual machine using Liberty dataset for various parsers across different dataset sizes, ranging from 1/1024 to 1/2 of the dataset.

### v2

@@ 4.4.1 Experiment depiction
LogLead [^6] https://github.com/EvoTestOps/LogLead is a tool designed to benchmark log representations and anomaly detection algorithms, such as log parsing results [mantyla2023loglead]. It features integrations with many public datasets and log parsers. LogLead is built with efficiency in mind, achieved by using the Polars [^7] https://github.com/pola-rs/polars library, a notably faster alternative to Pandas [^8] https://github.com/pandas-dev/pandas. As LogLead integrates existing datasets and parsers, it offers an environment to evaluate the computational efficiency of different log parsers over different datasets. Unlike the experiments in the previous sections, no dataset-specific or parser-specific optimizations were made, as they are not available on the LogLead platform.
The scripts we used for our LogLead experiments are available [^9] https://github.com/EvoTestOps/LogLead/tree/main/demo/parser_benchmark. For this experiment, we first loaded and preprocessed log data from nine different systems from past studies. This preprocessing involved loading data into Polars data frames and normalizing it using a standard regular expression in the LogLead Enhancer class. The datasets included BGL, Liberty, Spirit, and Thunderbird, originating from real-world high-performance computing usage [bgl]. The Hadoop[hadoopdataset] and HDFS[hdfs] datasets contain logs from experiments conducted on distributed file systems. Additionally, Trainticket [zhou2018fault] and Web-shop [^10] https://github.com/GoogleCloudPlatform/microservices-demo datasets represent experiment logs from modern microservice-based systems [yu2023nezha]. Finally, the Android2 dataset, not to be confused with the Android dataset, originates from long-lasting Android reliability testing runs [mantyla2022pinpointing].
@@
Parser implementations for AEL, Brain, and IpLom were taken from [loghubtools]. For Spell [^11] https://github.com/bave/pyspell and LenMa [^12] https://github.com/keiichishima/templateminer, we used the implementations available in the footnotes. For Drain, we used the so-called Drain3131313https://github.com/logpai/Drain3/ implementation. In the Drain configuration file, we turned off Drain’s internal log message normalization via regular expressions and disabled profiling to ensure maximal parsing speed. For other parsers, no such settings were available.
As our purpose was to investigate the benefits of parallelization, we decided to implement a paralyzation version of the IpLom parser. We named it PlIpLom, Pl being short for Polars. We followed the original IpLom paper and the Python-based source code implementation from [loghubtools], and we replaced every loop we encountered with native Polars operations if possible. Native Polars operations are known for their efficiency.
@@ 4.4.2 Experiment results
During our experiments, we set a run-time limit of 10 minutes (600 seconds for each parser on each dataset). Accordingly, if a run exceeded 10 minutes, we allowed it to finish, but no further runs were executed. The remaining results are marked as ”Out of Time Limit” (OTL) in the tables from VII to XI. We also mark crashed runs as ”Did Not Finish” (DNF). We conducted a re-run for all crashes; however, if the re-run also failed to complete, we marked it as DNF.
For small and medium datasets, we show only results from parsing full (1/1) datasets. For large data sets, we show fractions of the data as well. The tables showing the fraction results are in our replication package [^14] https://github.com/M3SOulu/tipping-detailed-result-tables/tree/main/4.4
@@
We have two small log datasets. The results for the Hadoop dataset (177,592 log lines) on the local machine (LM) and virtual machine (VM) are shown in Table VII. Tipping, Drain, and PlIplom are the three fastest parsers in both the VM and LM settings. Similar to Hadoop, TrainTicket is a small dataset (272,270 log lines), and its results are documented in Table VII. In the TrainTicket dataset, Drain is the fastest on the LM, followed by Tipping and Iplom. However, the performance ranking on the VM changes, with Tipping being the fastest, followed by Drain and Iplom.
TABLE VII: Parsing times (in seconds) for parsers across full of small data sets
Data Hadoop Hadoop Nezha-$T$ Nezha-$T$
Machine LM VM LM VM
AEL 6.41 13.29 16.86 37.60
Brain 9.81 21.18 32.58 63.20
Drain 0.93 2.22 1.74 3.70
Iplom 2.29 5.01 6.28 13.55
Lenma 29.54 86.78 122.89 349.02
Pliplom 1.66 2.80 9.69 14.72
Spell 47.52 105.34 78.96 164.07
Tipping 0.22 0.36 2.01 1.11
@@
Next, we have four medium-sized datasets. Web-shop, with 3,958,203 log lines, presents its results for the local machine (LM) and virtual machine (VM) in Table VIII. In the LM setting, the top three parsers are PlIplom, Tipping, and Drain, whereas in the VM, the order is Tipping, PlIplom, and Drain. Here, we begin to observe meaningful differences in parsing speed, with the difference between the first and third parsers approximately 9x in LM and 8x in VM.
@@
The results for the BGL dataset, containing 4,747,963 log lines, are displayed in Table VIII. In both computing environments, the top three parsers, in order, are Tipping, PlIplom, and Drain.
@@
The Android2 dataset, with 5,203,599 log lines, shows results in Table VIII. Tipping is the fastest parser in both the VM and LM, while the second and third places alternate between Drain and PlIplom in the VM and LM, respectively.
@@
Finally, the HDFS dataset, which consists of 11,175,629 log lines, is presented in Table VIII. In the LM, PlIplom is the fastest, followed by Tipping and Drain, while in the VM, Tipping leads, followed by PlIplom and Drain.
TABLE VIII: Parsing times (in seconds) for parsers across medium size datasets
Data Nezha-Shop Nezha-Shop BGL BGL Android2 Android2 HDFS HDFS
Machine LM VM LM VM LM VM LM VM
AEL 119.08 291.96 348.72 639.22 1657.02 OTL 338.08 727.55
Brain 191.24 357.70 169.00 354.50 210.54 468.88 DNF 1016.13
Drain 25.61 40.52 21.14 44.20 30.56 60.84 76.20 128.90
Iplom 39.52 106.90 43.58 100.47 57.86 119.40 119.42 256.24
Lenma OTL OTL OTL OTL OTL OTL OTL OTL
Pliplom 4.73 11.52 12.18 15.03 33.04 37.50 8.89 31.21
Spell 1556.72 OTL OTL OTL OTL OTL 821.56 OTL
Tipping 7.69 5.89 8.86 7.35 9.44 8.82 16.44 16.60
Finally, we have three large datasets executed only in the virtual machine (VM), as the local machine’s memory size was insufficient. In the VM, none of the parsing executions resulted in fully parsed datasets, as all of the parsers either exceeded the time limit or ran out of memory.
The Thunderbird dataset, with 211,212,192 log lines, shows parsing results in Table IX. The fastest parsers, in order, were Tipping, Iplom, and Drain. This order is consistent across the remaining two large datasets. The Liberty and Spirit datasets’ results are shown in Tables XI and X, respectively. Compared to its competitors, the only drawback noted for Tipping is its inability to parse half of the Liberty data, a task that PlIplom manages to accomplish.
@@
TABLE IX: Parsing times (in seconds) on the virtual machine using Thunderbird dataset for various parsers across different dataset sizes, ranging from 1/1024 to 1/2 of the dataset.
@@
TABLE X: Parsing times (in seconds) on the virtual machine using Spirit dataset for various parsers across different dataset sizes, ranging from 1/1024 to 1/4 of the dataset.
@@
TABLE XI: Parsing times (in seconds) on the virtual machine using Liberty dataset for various parsers across different dataset sizes, ranging from 1/1024 to 1/2 of the dataset.